# Optimized effective potential forces with the plane-wave and pseudopotential method

### v1

@@ I Introduction
Meta-GGAs, hybrid and MBPT-based functionals all have an explicit dependency on the KS orbitals, rather than the density. As a consequence, the xc potential, i.e., the functional derivative of the xc energy with respect to the density, does not have an analytical expression. Self-consistent calculations are, therefore, often performed within the generalised KS framework [12], or by allowing for larger variational freedom. In the case of hybrid functionals this implies the use of an integral operator, the nonlocal exchange potential [13], and, in the case of meta-GGAs, the use of a differential operator [14, 15]. The RPA functional has a natural extension within MBPT. Free variations with respect to the many-body Green’s function leads to a nonlocal and energy-dependent potential - the self-energy [16, 17, 18]. Due to the high complexity of such calculations there are only a few reported in the literature [19, 20, 21, 22].
@@
In order to remain within the KS formulation, the variational freedom has to be restricted such that the orbitals are generated by a local multiplicative potential. Functional differentiation with respect to the density via the KS orbitals leads to an integral equation, known as the optimized effective potential (OEP) equation, for determining the xc potential numerically [23, 24]. Tests on various systems have demonstrated that the OEP often gives total energies and densities in close agreement with results obtained within the generalised KS scheme [25, 26, 27]. On the other hand, except for the highest occupied molecular orbital (HOMO) eigenvalue [28], the spectrum will always be different. It is well known that the band gap of solids is strongly underestimated, even with the exact KS potential [29, 30]. In fact, one can show that the KS potential must jump with a constant when crossing integer particle numbers in the ensemble formulation of DFT [31]. Adding this constant, or the so-called derivative discontinuity correction, to the KS gap results in the true fundamental gap. Within hybrid functionals, the corrected KS gap is in good agreement with the gap obtained from the generalised KS scheme [32], and, within RPA, it becomes equivalent to the gap within theory [33, 34, 35].
@@
The numerical solution to the OEP equation has been the subject of numerous studies. A direct solution based on the inversion of the KS density response function is known to present numerical instabilities in Gaussian basis sets [36, 37, 38]. These issues are almost absent from OEP calculations resorting to plane-waves [39], real space grids [40, 41], or spline basis sets [18, 42, 43, 44, 45]. Furthermore, direct minimization [46] and iterative approaches [47, 48, 49] that circumvent the inversion of the KS density response function have been developed, making routine OEP calculations feasible on a wide range of systems. All-electron studies have focused on the role of the core-valence interaction and concluded that the pseudopotential approximation is valid also for OEP calculations [50, 41, 51, 52]. A review on orbital-dependent functionals and their numerical aspects can be found in Ref. 53.
@@ II Optimized effective potential
where is the initial trial potential. In each iteration, the coefficients are determined by minimizing the integral of and convergence is achieved when this value reaches a given threshold. The converged exchange potential is then used to update and the KS equation is solved non-self-consistently to generate a new set of orbitals that are used for solving the OEP equation another time. This cyclic procedure continues until the electronic density is found converged to a given threshold. An extension of this scheme to hybrid functionals and solids can be found in Refs. 32, 66.
@@ III Analytical forces with OEP
Following the notation in Ref. 67, we denote the set of nuclear positions as . At equilibrium geometry, the force exerted on each nucleus is zero. Calculating the second derivative of with respect to then yields interatomic force constants, which are necessary quantities for computing harmonic vibrational frequencies [68].
@@ III.2 Numerical test
In all our calculations, we used PBE optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [81] containing nonlocal projectors. The extra OEP force term in Eq. (24) is, therefore, necessary to include.
@@
We started by studying the water molecule H2O (see Fig. 3) using six different approximations: PBE, vdW-DF, HF, EXX, PBE025, and OEP25. All data have been obtained using a simulation cell of 25 Bohr, and results for 80 Ry and 200 Ry plane-wave cutoff are compared. The threshold on the energy convergence of PBE, vdW-DF, HF, and PBE025 was set to lowest possible value to ensure high accuracy of the numerical forces. The same was done for the two different thresholds used by EXX and OEP25 (see Sec. II). Regarding the atomic structure, the angle between the two O-H bonds has been fixed to 104.3°. Only the O-H(1) bond distance is varied while the length of O-H(2) is set to 0.97 Å. These parameters define a geometry for the water molecule that is close to the PBE equilibrium geometry. By performing several test calculations at different geometries, we found that the choice of starting geometry does not impact the results we obtained on H2O.
@@
In Fig. 3, the difference between analytical and numerical forces is calculated at a cutoff of 80 Ry. For every method, the scale of the error on FA is found to be relatively small. The worst agreement is in fact found with vdW-DF and PBE. The performance of the hybrid PBE025 functional is one order of magnitude better, and three orders of magnitude better with HF. If we now analyze the performance of the OEP methods, we notice that OEP25 produces a similar error to PBE025, which on average is about Ha/Å at 80 Ry. On the other hand, EXX fails to deliver the exceptionally small error seen with HF. While HF returns force differences of the order of Ha/Å, EXX results are closer in magnitude to PBE025 and OEP25. The difference observed between EXX and HF is probably related to the threshold used to solve the OEP equation. In Fig. 3 we present the same curves as in Fig. 3 but with a higher plane-wave cutoff of 200 Ry. An improvement of roughly one order of magnitude is seen in the error profile of all methods except HF, for which the error was already very small. The enhanced performance of the EXX approximation suggests that, for a given threshold used to solve the OEP equation, a higher plane-wave cutoff helps to improve the accuracy of the EXX potential. For the OEP methods, the extra source of error related to the evaluation of the local KS potential via the OEP equation seems to have a small impact. This is evident from the good agreement observed between OEP25 and PBE025 in Figs. 3 and 3. Only small irregularities in the error of the force can be seen with OEP25.
@@
To investigate whether the high accuracy we observe on molecules also extends to solids, we have performed the same analysis of forces on the -quartz phase of SiO2. For this solid that belongs to the P3221 space group, cell parameters and atomic positions have first been relaxed at the PBE level using a uniform 333 Monkhorst-Pack grid of k points and a cutoff of 80 Ry. On the optimized structure, one of the silicon ions has then been displaced away from its equilibrium position along the Cartesian x-axis, and the change in the force exerted on this ion has been monitored using PBE, vdW-DF, PBE025 and OEP25 at 80 Ry and 200 Ry. The data obtained are presented in Fig. 4.
@@
Similarly to the conclusions drawn from Fig. 3, we find on silica that PBE025 performs better than PBE, which in turn performs better than vdW-DF, at fixed planewave cutoff. The overall error on FA forces is, however, smaller on SiO2 as compared to H2O. This renders the irregularities in the error profile of OEP25 more apparent, despite their amplitudes being very small, about Ha/Å at 80 Ry. Given the magnitude of the force exerted on ion Si(1) in Fig. 4, the errors observed here are not expected to be of practical importance. As already noticed on H2O, using a larger cutoff leads to a decrease of the error in analytical forces for all methods. The OEP25 irregularities are also dampened, reducing from Ha/Å at 80 Ry to Ha/Å at 200 Ry. The different results obtained on SiO2 confirm a good accuracy of the OEP forces for periodic solids. We have been able to identify two main sources of errors in the analytical forces. The first error is method dependent. Different functionals may need different cutoffs to be fully converged. The second error is specific to the OEP method and is related to the accuracy of the local potential generated from the iterative solution of the OEP equation. Both of these errors can be reduced by increasing the plane-wave cutoff and by improving the convergency thresholds of the OEP and KS equations.
@@ IV Applications
The three systems investigated are the water molecule H2O, diamond, and the -quartz phase of SiO2. For H2O, we used the same functionals as tested in Sec. III, i.e. PBE, vdW-DF, HF, EXX, PBE025, and OEP25. For diamond and SiO2, only PBE, PBE025 and OEP25 have been used. For each functional we first relaxed the structure. The vibration modes were subsequently determined using the Phonopy Python package [84, 85]. Considering the atomic positions and the existing symmetries in the system, this code is able to generate several relevant supercell configurations through slight displacement of the elements. After computation of the analytical forces of each supercell, the existing vibration modes in the system can be predicted by Phonopy and their frequency calculated by finite difference of the forces using the central derivative formula.
@@ IV.2 Diamond
Diamond crystallizes in the Fd-3m space group. Thanks to the high symmetry, the atomic positions are fixed within the unit cell. Only the lattice parameter is left to vary in order to identify the optimized unit cell.
We have identified the suitable cell dimensions by monitoring the change of the total energy upon variations of the lattice parameter . We found that values of 6.700 Bohr for both OEP25 and PBE025, and 6.742 Bohr for PBE are optimal.
@@ V Conclusions and outlook
In this work, we have shown that OEP forces, within hybrid functionals, can be computed with a numerical accuracy similar to that obtained with commonly used functionals in DFT (e.g. PBE and vdW-DF). However, we also showed that special care is needed when employing nonlocal pseudopotentials. Since the OEP is based on a constrained optimization, an extra force term needs to be added to the standard Hellmann-Feynman expression for the forces. We implemented this term within the ACFDT package of the QUANTUM ESPRESSO distribution, that already computes the OEP potential. This allowed us to calculate forces, relax geometries, and determine phonon frequencies for a number of molecules and solids. Our different studies show that the local OEP exchange potential is a good approximation to the nonlocal exchange potential, being able to produce almost identical equilibrium structures and phonon frequencies.
@@ Acknowledgements.
The authors would like to thank Michele Casula and Lorenzo Paulatto for helpful discussions. The work was partially performed using HPC resources from GENCI-TGCC/CINES/IDRIS.References
- Hohenberg and Kohn [1964] P. Hohenberg and W. Kohn, Phys. Rev. 136, B864 (1964).
- Kohn and Sham [1965] W. Kohn and L. J. Sham, Phys. Rev. 140, A1133 (1965).
@@
- Fiolhais et al. [2003] C. Fiolhais, F. Nogueira, and M. Marques, A Primer in Density Functional Theory, 1st ed., Vol. 620 (Springer Berlin, Heidelberg, 2003).
- Kohn et al. [1996] W. Kohn, A. D. Becke, and R. G. Parr, The Journal of Physical Chemistry 100, 12974 (1996).
- Perdew and Schmidt [2001] J. P. Perdew and K. Schmidt, AIP Conference Proceedings 577, 1 (2001).
@@
- Perdew and Zunger [1981] J. P. Perdew and A. Zunger, Phys. Rev. B 23, 5048 (1981).
@@
- Becke [1993] A. D. Becke, The Journal of Chemical Physics 98, 5648 (1993).
@@
- Heyd et al. [2003] J. Heyd, G. E. Scuseria, and M. Ernzerhof, The Journal of Chemical Physics 118, 8207 (2003).
- Ren et al. [2012] X. Ren, P. Rinke, C. Joas, and M. Scheffler, J Mater Sci 47, 7447 (2012).
@@
- Seidl et al. [1996] A. Seidl, A. Görling, P. Vogl, J. A. Majewski, and M. Levy, Phys. Rev. B 53, 3764 (1996).
@@
- Paier et al. [2006] J. Paier, M. Marsman, K. Hummer, G. Kresse, I. C. Gerber, and J. G. Ángyán, The Journal of Chemical Physics 124, 154709 (2006).
- Eich and Hellgren [2014] F. G. Eich and M. Hellgren, The Journal of Chemical Physics 141, 224107 (2014).
- Yang et al. [2016] Z. H. Yang, H. Peng, J. Sun, and J. P. Perdew, Phys. Rev. B 93, 205205 (2016).
- Klein [1961] A. Klein, Phys. Rev. 121, 950 (1961).
- Dahlen et al. [2006] N. E. Dahlen, R. van Leeuwen, and U. von Barth, Phys. Rev. A 73, 012511 (2006).
- Hellgren and von Barth [2007] M. Hellgren and U. von Barth, Phys. Rev. B 76, 075107 (2007).
- Holm and von Barth [1998] B. Holm and U. von Barth, Phys. Rev. B 57, 2108 (1998).
- Stan et al. [2006] A. Stan, N. E. Dahlen, and R. Van Leeuwen, Europhysics Letters 76, 298 (2006).
- Caruso et al. [2012] F. Caruso, P. Rinke, X. Ren, M. Scheffler, and A. Rubio, Phys. Rev. B 86, 081102(R) (2012).
@@
- Grumet et al. [2018] M. Grumet, P. Liu, M. Kaltak, J. c. v. Klimeš, and G. Kresse, Phys. Rev. B 98, 155143 (2018).
- Sharp and Horton [1953] R. T. Sharp and G. K. Horton, Phys. Rev. 90, 317 (1953).
- Talman and Shadwick [1976] J. D. Talman and W. F. Shadwick, Phys. Rev. A 14, 36 (1976).
- Aashamar et al. [1979] K. Aashamar, T. M. Luke, and J. D. Talman, Phys. Rev. A 19, 6 (1979).
- Krieger et al. [1992a] J. B. Krieger, Y. Li, and G. J. Iafrate, Phys. Rev. A 45, 101 (1992a).
- Kim et al. [2017] J. Kim, K. Hong, S.-Y. Hwang, S. Ryu, S. Choi, and W. Y. Kim, Phys. Chem. Chem. Phys. 19, 10177 (2017).
- Almbladh and von Barth [1985] C.-O. Almbladh and U. von Barth, Phys. Rev. B 31, 3231 (1985).
- Li et al. [1991] Y. Li, J. B. Krieger, M. R. Norman, and G. J. Iafrate, Phys. Rev. B 44, 10437 (1991).
@@
- Perdew [1985] J. P. Perdew, International Journal of Quantum Chemistry 28, 497 (1985).
@@
- Perdew et al. [1982] J. P. Perdew, R. G. Parr, M. Levy, and J. L. Balduz, Phys. Rev. Lett. 49, 1691 (1982).
@@
- Hellgren et al. [2021] M. Hellgren, L. Baguet, M. Calandra, F. Mauri, and L. Wirtz, Phys. Rev. B 103, 075101 (2021).
- Niquet and Gonze [2004] Y. M. Niquet and X. Gonze, Phys. Rev. B 70, 245115 (2004).
@@
- Grüning et al. [2006] M. Grüning, A. Marini, and A. Rubio, Phys. Rev. B 74, 161103(R) (2006).
- Klimeš and Kresse [2014] J. Klimeš and G. Kresse, The Journal of Chemical Physics 140, 054516 (2014).
- Görling [1999] A. Görling, Phys. Rev. Lett. 83, 5459 (1999).
- Heßelmann et al. [2007] A. Heßelmann, A. W. Götz, F. Della Sala, and A. Görling, The Journal of Chemical Physics 127, 054102 (2007).
- Trushin and Görling [2021] E. Trushin and A. Görling, The Journal of Chemical Physics 155, 054109 (2021).
@@
- Städele et al. [1999] M. Städele, M. Moukara, J. A. Majewski, P. Vogl, and A. Görling, Phys. Rev. B 59, 10031 (1999).
- Jiang and Engel [2005] H. Jiang and E. Engel, The Journal of Chemical Physics 123, 224102 (2005).

### v2

@@ I Introduction
Meta-GGAs, hybrid and MBPT-based functionals all have an explicit dependency on the KS orbitals, rather than the density. As a consequence, the xc potential, i.e., the functional derivative of the xc energy with respect to the density, does not have an analytical expression. Self-consistent calculations are, therefore, often performed within the generalised KS framework [12, 13, 14], or by allowing for larger variational freedom. In the case of hybrid functionals this implies the use of an integral operator, the nonlocal exchange potential [15], and, in the case of meta-GGAs, the use of a differential operator [16, 17]. The RPA functional has a natural extension within MBPT. Free variations with respect to the many-body Green’s function leads to a nonlocal and energy-dependent potential - the self-energy [18, 19, 20]. Due to the high complexity of such calculations there are only a few reported in the literature [21, 22, 23, 24].
@@
In order to remain within the KS formulation, the variational freedom has to be restricted such that the orbitals are generated by a local multiplicative potential. Functional differentiation with respect to the density via the KS orbitals leads to an integral equation, known as the optimized effective potential (OEP) equation, for determining the xc potential numerically [25, 26]. Tests on various systems have demonstrated that the OEP often gives total energies and densities in close agreement with results obtained within the generalised KS scheme [27, 28, 29]. On the other hand, except for the highest occupied molecular orbital (HOMO) eigenvalue [30], the spectrum will always be different. The KS virtual orbitals can, for example, be shown to provide a better description of the optical excitation energies [31], but the fundamental gap is strongly underestimated, even with the exact KS potential [32, 33]. One can show that the KS potential must jump with a constant when crossing integer particle numbers in the ensemble formulation of DFT [34]. Adding this constant, or the so-called derivative discontinuity correction, to the KS gap results in the true fundamental gap. Within hybrid functionals, the corrected KS gap is in good agreement with the gap obtained from the generalised KS scheme [35], and same is expected to be true for meta-GGAs [14]. Within RPA, it becomes equivalent to the gap within theory [36, 37, 38].
@@
The numerical solution to the OEP equation has been the subject of numerous studies. A direct solution based on the inversion of the KS density response function is known to present numerical instabilities in Gaussian basis sets [39, 40, 41]. These issues are almost absent from OEP calculations resorting to plane-waves [42], real space grids [43, 44], or spline basis sets [20, 45, 46, 47, 48]. Furthermore, direct minimization [49] and iterative approaches [50, 51, 52] that circumvent the inversion of the KS density response function have been developed, making routine OEP calculations feasible on a wide range of systems. All-electron studies have focused on the role of the core-valence interaction and concluded that the pseudopotential approximation is valid also for OEP calculations [53, 54, 44, 55, 56]. A review on orbital-dependent functionals and their numerical aspects can be found in Ref. 57.
@@ II Optimized effective potential
where is the initial trial potential. In each iteration, the coefficients are determined by minimizing the integral of and convergence is achieved when this value reaches a given threshold. The converged exchange potential is then used to update and the KS equation is solved non-self-consistently to generate a new set of orbitals that are used for solving the OEP equation another time. This cyclic procedure continues until the electronic density is found converged to a given threshold. An extension of this scheme to hybrid functionals and solids can be found in Refs. 35, 69.
@@ III Analytical forces with OEP
Following the notation in Ref. 70, we denote the set of nuclear positions as . At equilibrium geometry, the force exerted on each nucleus is zero. Calculating the second derivative of with respect to then yields interatomic force constants, which are necessary quantities for computing harmonic vibrational frequencies [71].
@@ III.2 Numerical test
In principle, one should use pseudopotentials optimized for the specific functional used [53]. However, since our objectives are to study the numerical precision and to compare OEP to generalized KS calculations we have settled for PBE optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [84]. Independently of the functional used to optimize the pseudopotential they will, in general, contain nonlocal projectors and the extra OEP force term in Eq. (24) is necessary to include.
@@
We started by studying the water molecule H2O (see Fig. 3) using five different approximations: PBE, HF, EXX, PBE025, and OEP25. All data have been obtained using a simulation cell of 25 Bohr, and results for 80 Ry and 200 Ry plane-wave cutoff are compared. The threshold on the energy convergence of PBE, HF, and PBE025 was set to lowest possible value to ensure high accuracy of the numerical forces. The same was done for the two different thresholds used by EXX and OEP25 (see Sec. II). Regarding the atomic structure, the angle between the two O-H bonds has been fixed to 104.3°. Only the O-H(1) bond distance is varied while the length of O-H(2) is set to 0.97 Å. These parameters define a geometry for the water molecule that is close to the PBE equilibrium geometry. By performing several test calculations at different geometries, we found that the choice of starting geometry does not impact the results we obtained on H2O.
@@
In Fig. 3, the difference between analytical and numerical forces is calculated at a cutoff of 80 Ry. For every method, the scale of the error on FA is found to be relatively small. The performance of the hybrid PBE025 functional is one order of magnitude better than PBE, and three orders of magnitude better with HF. If we now analyze the performance of the OEP methods, we notice that OEP25 produces a similar error to PBE025, which on average is about Ha/Å at 80 Ry. On the other hand, EXX fails to deliver the exceptionally small error seen with HF. While HF returns force differences of the order of Ha/Å, EXX results are closer in magnitude to PBE025 and OEP25. The difference observed between EXX and HF is probably related to the threshold used to solve the OEP equation. In Fig. 3 we present the same curves as in Fig. 3 but with a higher plane-wave cutoff of 200 Ry. An improvement of roughly one order of magnitude is seen in the error profile of all methods except HF, for which the error was already very small. The enhanced performance of the EXX approximation suggests that, for a given threshold used to solve the OEP equation, a higher plane-wave cutoff helps to improve the accuracy of the EXX potential. For the OEP methods, the extra source of error related to the evaluation of the local KS potential via the OEP equation seems to have a small impact. This is evident from the good agreement observed between OEP25 and PBE025 in Figs. 3 and 3. Only small irregularities in the error of the force can be seen with OEP25.
@@
To investigate whether the high accuracy we observe on molecules also extends to solids, we have performed the same analysis of forces on the -quartz phase of SiO2. For this solid that belongs to the P3221 space group, cell parameters and atomic positions have first been relaxed at the PBE level using a uniform 333 Monkhorst-Pack grid of k points and a cutoff of 80 Ry. On the optimized structure, one of the silicon ions has then been displaced away from its equilibrium position along the Cartesian x-axis, and the change in the force exerted on this ion has been monitored using PBE, PBE025 and OEP25 at 80 Ry and 200 Ry. The data obtained are presented in Fig. 4.
@@
Similarly to the conclusions drawn from Fig. 3, we find on silica that PBE025 performs better than PBE at fixed planewave cutoff. The overall error on FA forces is, however, smaller on SiO2 as compared to H2O. This renders the irregularities in the error profile of OEP25 more apparent, despite their amplitudes being very small, about Ha/Å at 80 Ry. Given the magnitude of the force exerted on ion Si(1) in Fig. 4, the errors observed here are not expected to be of practical importance. As already noticed on H2O, using a larger cutoff leads to a decrease of the error in analytical forces for all methods. The OEP25 irregularities are also dampened, reducing from Ha/Å at 80 Ry to Ha/Å at 200 Ry. The different results obtained on SiO2 confirm a good accuracy of the OEP forces for periodic solids. We have been able to identify two main sources of errors in the analytical forces. The first error is method dependent. Different functionals may need different cutoffs to be fully converged. The second error is specific to the OEP method and is related to the accuracy of the local potential generated from the iterative solution of the OEP equation. Both of these errors can be reduced by increasing the plane-wave cutoff and by improving the convergency thresholds of the OEP and KS equations.
@@ IV Applications
The three systems investigated are the water molecule H2O, diamond, and the -quartz phase of SiO2. For H2O, we used the same functionals as tested in Sec. III, i.e. PBE, HF, EXX, PBE025, and OEP25. For diamond and SiO2, only PBE, PBE025 and OEP25 have been used. For each functional we first relaxed the structure. The vibration modes were subsequently determined using the Phonopy Python package [87, 88]. Considering the atomic positions and the existing symmetries in the system, this code is able to generate several relevant supercell configurations through slight displacement of the elements. After computation of the analytical forces of each supercell, the existing vibration modes in the system can be predicted by Phonopy and their frequency calculated by finite difference of the forces using the central derivative formula.
@@ IV.2 Diamond
Diamond crystallizes in the Fd-3m space group. Thanks to the high symmetry, the atomic positions are fixed within the unit cell. Only the lattice parameter is left to vary in order to identify the optimized unit cell. We have identified the suitable cell dimensions by monitoring the change of the total energy upon variations of the lattice parameter . The values of 6.700 Bohr for both OEP25 and PBE025, and 6.742 Bohr for PBE are optimal.
We first calculated the electronic band structure along the L - - X high symmetry path within PBE025 and OEP25 (see Fig. 5). The OEP25 bands are easy to generate using post-processing tools since the local exchange potential is -independent. For PBE025 we only have the band energies on the -point grid used in the self-consistent calculation. The bands have, therefore, been spline-interpolated between these points. The band structures are clearly different in the two approximations. The first conduction band is shifted by approximately 2 eV in PBE025. This difference is expected and related to the derivative discontinuity within the OEP method [32, 33]. Adding the derivative discontinuity correction to the OEP25 result returns a gap in very good agreement with PBE025, with an energy difference of 0.02 eV.
@@ V Conclusions and outlook
In this work, we have shown that OEP forces, within hybrid functionals, can be computed with a numerical accuracy similar to that obtained with commonly used functionals in DFT. However, we also showed that special care is needed when employing nonlocal pseudopotentials. Since the OEP is based on a constrained optimization, an extra force term needs to be added to the standard Hellmann-Feynman expression for the forces. We implemented this term within the ACFDT package of the QUANTUM ESPRESSO distribution, that already computes the OEP potential. This allowed us to calculate forces, relax geometries, and determine phonon frequencies for a number of molecules and solids. Our different studies show that the local OEP exchange potential is a good approximation to the nonlocal exchange potential, being able to produce almost identical equilibrium structures and phonon frequencies.
@@ Acknowledgements.
The authors would like to thank Michele Casula and Lorenzo Paulatto for helpful discussions. The work was partially performed using HPC resources from GENCI-TGCC/CINES/IDRIS (Grant No. A0150914650).References
- Hohenberg and Kohn [1964] P. Hohenberg and W. Kohn, Inhomogeneous Electron Gas, Phys. Rev. 136, B864–B871 (1964).
- Kohn and Sham [1965] W. Kohn and L. J. Sham, Self-Consistent Equations Including Exchange and Correlation Effects, Phys. Rev. 140, A1133–A1138 (1965).
@@
- Fiolhais et al. [2003] C. Fiolhais, F. Nogueira, and M. A. L. Marques, A Primer in Density Functional Theory, 1st ed., Vol. 620 (Springer Berlin, Heidelberg, 2003).
- Kohn et al. [1996] W. Kohn, A. D. Becke, and R. G. Parr, Density Functional Theory of Electronic Structure, The Journal of Physical Chemistry 100, 12974-12980 (1996).
- Perdew and Schmidt [2001] J. P. Perdew and K. Schmidt, Jacob’s ladder of density functional approximations for the exchange-correlation energy, AIP Conference Proceedings 577, 1-20 (2001).
@@
- Perdew and Zunger [1981] J. P. Perdew and A. Zunger, Self-interaction correction to density-functional approximations for many-electron systems, Phys. Rev. B 23, 5048–5079 (1981).
@@
- Becke [1993] A. D. Becke, Density‐functional thermochemistry. III. The role of exact exchange, The Journal of Chemical Physics 98, 5648-5652 (1993).
@@
- Heyd et al. [2003] J. Heyd, G. E. Scuseria, and M. Ernzerhof, Hybrid functionals based on a screened Coulomb potential, The Journal of Chemical Physics 118, 8207-8215 (2003).
- Ren et al. [2012] X. Ren, P. Rinke, C. Joas, and M. Scheffler, Random-phase approximation and its applications in computational chemistry and materials science, Journal of Materials Science 47, 7447-7471 (2012).
@@
- Seidl et al. [1996] A. Seidl, A. Görling, P. Vogl, J. A. Majewski, and M. Levy, Generalized Kohn-Sham schemes and the band-gap problem, Phys. Rev. B 53, 3764–3774 (1996).
- Garrick et al. [2020] R. Garrick, A. Natan, T. Gould, and L. Kronik, Exact Generalized Kohn-Sham Theory for Hybrid Functionals, Phys. Rev. X 10, 021040 (2020).
- Lebeda et al. [2023] T. Lebeda, T. Aschebrock, J. Sun, L. Leppert, and S. Kümmel, Right band gaps for the right reason at low computational cost with a meta-GGA, Phys. Rev. Mater. 7, 093803 (2023).
@@
- Paier et al. [2006] J. Paier, M. Marsman, K. Hummer, G. Kresse, I. C. Gerber, and J. G. Ángyán, Screened hybrid density functionals applied to solids, The Journal of Chemical Physics 124, 154709 (2006).
- Eich and Hellgren [2014] F. G. Eich and M. Hellgren, Derivative discontinuity and exchange-correlation potential of meta-GGAs in density-functional theory, The Journal of Chemical Physics 141, 224107 (2014).
- Yang et al. [2016] Z. Yang, H. Peng, J. Sun, and J. P. Perdew, More realistic band gaps from meta-generalized gradient approximations: Only in a generalized Kohn-Sham scheme, Phys. Rev. B 93, 205205 (2016).
- Klein [1961] A. Klein, Perturbation Theory for an Infinite Medium of Fermions. II, Phys. Rev. 121, 950–956 (1961).
- Dahlen et al. [2006] N. E. Dahlen, R. van Leeuwen, and U. von Barth, Variational energy functionals of the Green function and of the density tested on molecules, Phys. Rev. A 73, 012511 (2006).
- Hellgren and von Barth [2007] M. Hellgren and U. von Barth, Correlation potential in density functional theory at the GWA level: Spherical atoms, Phys. Rev. B 76, 075107 (2007).
- Holm and von Barth [1998] B. Holm and U. von Barth, Fully self-consistent self-energy of the electron gas, Phys. Rev. B 57, 2108–2117 (1998).
- Stan et al. [2006] A. Stan, N. E. Dahlen, and R. Van Leeuwen, Fully self-consistent calculations for atoms and molecules, Europhys. Lett. 76, 298–304 (2006).
- Caruso et al. [2012] F. Caruso, P. Rinke, X. Ren, M. Scheffler, and A. Rubio, Unified description of ground and excited states of finite systems: The self-consistent approach, Phys. Rev. B 86, 081102(R) (2012).
@@
- Grumet et al. [2018] M. Grumet, P. Liu, M. Kaltak, J. c. v. Klimeš, and G. Kresse, Beyond the quasiparticle approximation: Fully self-consistent calculations, Phys. Rev. B 98, 155143 (2018).
- Sharp and Horton [1953] R. T. Sharp and G. K. Horton, A Variational Approach to the Unipotential Many-Electron Problem, Phys. Rev. 90, 317–317 (1953).
- Talman and Shadwick [1976] J. D. Talman and W. F. Shadwick, Optimized effective atomic central potential, Phys. Rev. A 14, 36–40 (1976).
- Aashamar et al. [1979] K. Aashamar, T. M. Luke, and J. D. Talman, Properties of single-term atomic states calculated in a variationally optimized-local-central-potential model, Phys. Rev. A 19, 6–16 (1979).
- Krieger et al. [1992a] J. B. Krieger, Y. Li, and G. J. Iafrate, Construction and application of an accurate local spin-polarized Kohn-Sham potential with integer discontinuity: Exchange-only theory, Phys. Rev. A 45, 101–126 (1992a).
- Kim et al. [2017] J. Kim, K. Hong, S.-Y. Hwang, S. Ryu, S. Choi, and W. Y. Kim, Effects of the locality of a potential derived from hybrid density functionals on Kohn–Sham orbitals and excited states, Phys. Chem. Chem. Phys. 19, 10177-10186 (2017).
- Almbladh and von Barth [1985] C.-O. Almbladh and U. von Barth, Exact results for the charge and spin densities, exchange-correlation potentials, and density-functional eigenvalues, Phys. Rev. B 31, 3231–3244 (1985).
- Baerends et al. [2013] E. J. Baerends, O. V. Gritsenko, and R. van Meer, The Kohn–Sham gap, the fundamental gap and the optical gap: the physical meaning of occupied and virtual Kohn–Sham orbital energies, Phys. Chem. Chem. Phys. 15, 16408-16425 (2013).
- Li et al. [1991] Y. Li, J. B. Krieger, M. R. Norman, and G. J. Iafrate, Band-structure calculations of noble-gas and alkali halide solids using accurate Kohn-Sham potentials with self-interaction correction, Phys. Rev. B 44, 10437–10443 (1991).
@@
- Perdew [1985] J. P. Perdew, Density functional theory and the band gap problem, International Journal of Quantum Chemistry 28, 497-523 (1985).
@@
- Perdew et al. [1982] J. P. Perdew, R. G. Parr, M. Levy, and J. L. Balduz, Density-Functional Theory for Fractional Particle Number: Derivative Discontinuities of the Energy, Phys. Rev. Lett. 49, 1691–1694 (1982).
@@
- Hellgren et al. [2021] M. Hellgren, L. Baguet, M. Calandra, F. Mauri, and L. Wirtz, Electronic structure of from a quasi-self-consistent approach, Phys. Rev. B 103, 075101 (2021).
- Niquet and Gonze [2004] Y. M. Niquet and X. Gonze, Band-gap energy in the random-phase approximation to density-functional theory, Phys. Rev. B 70, 245115 (2004).
@@
- Grüning et al. [2006] M. Grüning, A. Marini, and A. Rubio, Effect of spatial nonlocality on the density functional band gap, Phys. Rev. B 74, 161103(R) (2006).
- Klimeš and Kresse [2014] J. Klimeš and G. Kresse, Kohn-Sham band gaps and potentials of solids from the optimised effective potential method within the random phase approximation, The Journal of Chemical Physics 140, 054516 (2014).
- Görling [1999] A. Görling, New KS Method for Molecules Based on an Exchange Charge Density Generating the Exact Local KS Exchange Potential, Phys. Rev. Lett. 83, 5459–5462 (1999).
- Heßelmann et al. [2007] A. Heßelmann, A. W. Götz, F. Della Sala, and A. Görling, Numerically stable optimized effective potential method with balanced Gaussian basis sets, The Journal of Chemical Physics 127, 054102 (2007).
- Trushin and Görling [2021] E. Trushin and A. Görling, Numerically stable optimized effective potential method with standard Gaussian basis sets, The Journal of Chemical Physics 155, 054109 (2021).
@@
- Städele et al. [1999] M. Städele, M. Moukara, J. A. Majewski, P. Vogl, and A. Görling, Exact exchange Kohn-Sham formalism applied to semiconductors, Phys. Rev. B 59, 10031–10043 (1999).
- Jiang and Engel [2005] H. Jiang and E. Engel, Second-order Kohn-Sham perturbation theory: Correlation potential for atoms in a cavity, The Journal of Chemical Physics 123, 224102 (2005).